OpenStudy (anonymous):

Find the equation, in slope- intercept form, for the line parallel to the given line and passing through the indicated point.

OpenStudy (anonymous):

Parallel to 6x+2y=19 and passing through (-6,-13)

OpenStudy (anonymous):

what is the slope of 6x+2y=19? and then what is the line with that slope passing through (-6,-13)? and we want it in y=mx+b form

OpenStudy (apoorvk):

okay now the line is suppose ax+by=c now for the new line, 'a' and 'b' will be the same as its parallel to the first one. but 'c' value will be different. say its 'k' for the new line. so the new line becomes: ax+by=k. now put in the value of (x,y) from the coordinates of the point through which this new line passes, and get the value of 'k'. you will have your reqd. equation!! post the equation you get and verify!!

OpenStudy (anonymous):

Yes that is correct philips.

OpenStudy (anonymous):

Can you help me solve it philips13?

OpenStudy (anonymous):

Sorry i am terrible at math and i am not sure how to solve it :/

OpenStudy (apoorvk):